# On local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms

### v1

@@ Abstract.
In this article, we generalize some results in Chan-Yuan (2025) to local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms. In particular, we obtain further rigidity theorems and non-existence theorems for such maps.
@@ 1. Introduction
In [Ca53], Calabi has proved that there does not exist a local holomorphic isometric embedding between complex space forms of different types. This motivates our study in Chan-Yuan [CY25], and we have partially generalized Calabi’s result to the case of local holomorphic maps preserving $(p,p)$-forms (cf. Chan-Yuan [CY25, Theorem 1.4]). Recently, Arezzo-Li-Loi [ALL25, Corollary 1.7] has proved that there does not exist a Ricci-flat non-compact Kähler submanifold of any finite dim. complex projective space equipped with the Fubini–Study metric. Hence, it follows from Chan-Yuan [CY25, Theorem 1.4 and Remark 3.9] and Arezzo-Li-Loi [ALL25, Corollary 1.7] that if $(M,\omega_{M})$ and $(N,\omega_{N})$ are complex space forms of complex dimensions $m$ and $n$ respectively such that $M$ and $N$ are of different types, then for any real constant $\lambda>0$ and any integer $p$, $1\leq p\leq\min\{m,n\}$, there does not exist a local holomorphic map $F:(M;x_{0})\to(N;F(x_{0}))$ such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ holds. This yields a complete generalization of Calabi’s result. (Noting that in Chan-Yuan [CY25, Theorem 1.4] we exclude the case where the universal cover of $M$ is biholomorphic to $\mathbb{C}^{m}$, $N=\mathbb{P}^{n}$ and $2\leq p=m<n$.)
@@ 4.2. The Umehara algebra on bounded symmetric domains and its applications
for some homogeneous polynomials $G^{j}_{l}$, $1\leq l\leq N^{\prime}_{j}$, of degree $\geq 2$ (cf. Chan-Xiao-Yuan [CXY17, Section 4]). We compute
$\widetilde{\phi}_{m}(\zeta)=\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{m_{1}-1}&&&\\ &&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&\ddots&\\ &&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{m_{1}+1}},$
which does not lie in $\Lambda(\Delta)$ by [Um88], a plain contradiction. Hence, $\phi_{m}\not\in\Lambda(D)$. ∎
@@
for $1\leq l_{1}<\cdots<l_{p}\leq m$ and $1\leq k_{1}<\cdots<k_{p}\leq m$, where $U^{\prime}$ is a simply connected open neighborhood of $\bf 0$} in $U$ and $\lVert F^{j}(w)\rVert$ denotes the standard complex Euclidean norm of $F^{j}(w)$ in $\mathbb{C}^{N_{j}}$. Let $U^{\prime\prime}$ be a simply connected open neighborhood of $\bf 0$} in $\{\zeta\in\mathbb{C}:(\zeta,{\bf 0})\in U^{\prime}\}$. Now, we put $p=m$. By restricting to the minimal disk $\Delta_{0}:=D\cap\{(w_{1},\ldots,w_{m})\in\mathbb{C}^{m}:w_{j}=0\text{ for }j=2,\ldots,m\}$, we have $\phi_{m}(\zeta,{\bf 0})={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{m_{1}+1}}$ from the proof of Lemma 4.3, and thus by Chan-Yuan [CY25, Proof of Proposition 3.7] we have
$\psi_{m}(\zeta):={\prod_{j=1}^{l}(1+\lVert F^{j}(\zeta,{\bf 0})\rVert^{2})^{2m}\over(1-|\zeta|^{2})^{m_{1}+1}}\in\Lambda(U^{\prime\prime}).$

### v2

@@ Abstract.
In this article, we generalize some results in Chan-Yuan [Ann. Sc. Norm. Super. Pisa Cl. Sci. (5) 26 (2025), 619–644] to local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms. In particular, we obtain further rigidity theorems and non-existence theorems for such maps.
@@ 1. Introduction
In [Ca53], Calabi has proved that there does not exist a local holomorphic isometric embedding between complex space forms of different types. This motivates our study in Chan-Yuan [CY25], and we have partially generalized Calabi’s result to the case of local holomorphic maps preserving $(p,p)$-forms (cf. Chan-Yuan [CY25, Theorem 1.4]). Recently, Arezzo-Li-Loi [ALL25, Corollary 1.7] have proved that there does not exist a Ricci-flat non-compact Kähler submanifold of any finite dim. complex projective space equipped with the Fubini–Study metric. Hence, it follows from Chan-Yuan [CY25, Theorem 1.4 and Remark 3.9] and Arezzo-Li-Loi [ALL25, Corollary 1.7] that if $(M,\omega_{M})$ and $(N,\omega_{N})$ are complex space forms of complex dimensions $m$ and $n$ respectively such that $M$ and $N$ are of different types, then for any real constant $\lambda>0$ and any integer $p$, $1\leq p\leq\min\{m,n\}$, there does not exist a local holomorphic map $F:(M;x_{0})\to(N;F(x_{0}))$ such that $F^{*}\omega_{N}^{p}=\lambda\omega_{M}^{p}$ holds. This yields a complete generalization of Calabi’s result. (Noting that in Chan-Yuan [CY25, Theorem 1.4] we exclude the case where the universal cover of $M$ is biholomorphic to $\mathbb{C}^{m}$, $N=\mathbb{P}^{n}$ and $2\leq p=m<n$.)
@@ 4.2. The Umehara algebra on bounded symmetric domains and its applications
for some homogeneous polynomials $G^{j}_{l}$, $1\leq l\leq N^{\prime}_{j}$, of degree $\geq 2$ (cf. Chan-Xiao-Yuan [CXY17, Section 4]). By using Mok [Mo14, Lemma 3 and Lemma 5], we compute
$\begin{split}\widetilde{\phi}_{m}(\zeta)=&\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{p_{1}}&&&&\\ &&\lambda_{1}{\bf I}_{m_{1}-p_{1}-1}&&&\\ &&&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&&\ddots&\\ &&&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}\\ =&{\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}},\end{split}$
which does not lie in $\Lambda(\Delta)$ by Umehara [Um88], a plain contradiction. Here, $p_{1}$ is a non-negative integer depending only on $D_{1}$. Hence, $\phi_{m}\not\in\Lambda(D)$. ∎
@@
for $1\leq l_{1}<\cdots<l_{p}\leq m$ and $1\leq k_{1}<\cdots<k_{p}\leq m$, where $U^{\prime}$ is a simply connected open neighborhood of $\bf 0$} in $U$ and $\lVert F^{j}(w)\rVert$ denotes the standard complex Euclidean norm of $F^{j}(w)$ in $\mathbb{C}^{N_{j}}$. Let $U^{\prime\prime}$ be a simply connected open neighborhood of $\bf 0$} in $\{\zeta\in\mathbb{C}:(\zeta,{\bf 0})\in U^{\prime}\}$. Now, we put $p=m$. By restricting to the minimal disk $\Delta_{0}:=D\cap\{(w_{1},\ldots,w_{m})\in\mathbb{C}^{m}:w_{j}=0\text{ for }j=2,\ldots,m\}$, we have $\phi_{m}(\zeta,{\bf 0})={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}}$ from the proof of Lemma 4.3, and thus by Chan-Yuan [CY25, Proof of Proposition 3.7] we have
$\psi_{m}(\zeta):={\prod_{j=1}^{l}(1+\lVert F^{j}(\zeta,{\bf 0})\rVert^{2})^{2m}\over(1-|\zeta|^{2})^{p_{1}+2}}\in\Lambda(U^{\prime\prime}).$